OpenStudy (anonymous):

MEDALS AND FAN WITHIN FEW MINUTES Carlisle conducted an experiment to determine if the there is a difference in mean body temperature for men and women. He found that the mean body temperature for a sample of 100 men was 97.9 with a population standard deviation of 0.57 and the mean body temperature for a sample of 100 women was 98.6 with a population standard deviation of 0.55. Assuming the population of body temperatures for men and women is normally distributed, calculate the 99% confidence interval and the margin of error for the mean body temperature for both men and women. Using compl

OpenStudy (anonymous):

really not looking to get all wordy, someone just help me get the answer please

OpenStudy (kropot72):

The formula that you need for the confidence intervals is: \[\large \bar {x}\pm2.576\frac{\sigma}{\sqrt{n}}\]

OpenStudy (kropot72):

The margin of error is given by: \[\large 2.576\frac{\sigma}{\sqrt{n}}\]